OpenStudy (anonymous):

What is the name given to the difference between the minimum and maximum values for a data set? a. interquartile range b.median c. quartile d. range

OpenStudy (sriram):

interquartile range i guess

OpenStudy (zarkon):

d

OpenStudy (zarkon):

the interquartile range is the difference between the 3rd and 1st quartile IQR=Q3-Q1

OpenStudy (sriram):

but then the range is like (1,3) then diff= 3-1=2 so range and difference are not the same

OpenStudy (zarkon):

Range is defined to be the the largest data value minus the smallest data value